OpenStudy (anonymous):

If the ratio of the radii of two spheres is 8:9, what is the ratio of the volumes of the two spheres?

OpenStudy (anonymous):

What is the formula for a volume of a sphere?

OpenStudy (anonymous):

Do you know?

OpenStudy (anonymous):

\[8.9^3=704.969 \]

OpenStudy (amistre64):

elias, someone else is trying to offer the asker guidance, please do not come in and snipe the question

OpenStudy (anonymous):

The volume of sphere is \[ 4 \pi \frac {r^3} 3 \]

OpenStudy (amistre64):

http://openstudy.com/code-of-conduct